OpenStudy (anonymous):

solve: 0.25 = log16^x

OpenStudy (anonymous):

\[\log _{a}b =y\] is the same as writing \[a ^{y} =b\]

OpenStudy (anonymous):

a=10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now its 10^(0.25) = 16^x

OpenStudy (anonymous):

yes...that is it

OpenStudy (anonymous):

now what hahah

OpenStudy (raden):

log16^x = x log 16

OpenStudy (anonymous):

so play the 10^(0.25) on a calculator

OpenStudy (anonymous):

right, thats 1.77827

OpenStudy (anonymous):

from there, you can do the reverse of what i first wrote.......realise what you have is in the form \[a ^{y} = b\]

OpenStudy (anonymous):

so you can rewrite in the log form to find x..

OpenStudy (anonymous):

do you understand what i am saying???

OpenStudy (anonymous):

kind of except i dont know what to write..

OpenStudy (lasttccasey):

.25 = log(2^(4x)) using log rules you can bring the 4x out front: .25 = 4xlog(2) log(2) is a number so:\[\frac{ .25 }{ 4*\log(2) } = x\]

OpenStudy (anonymous):